OpenStudy (anonymous):

x^2+y^2+22x-14y+49=0 Parabola, Circle, Ellipse, or Hyperbola?

OpenStudy (anonymous):

compleate the squares first

OpenStudy (anonymous):

how do i do that? O:

OpenStudy (anonymous):

which level you are studying?

OpenStudy (anonymous):

err 4? o_o

OpenStudy (anonymous):

is that highschool?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so compleating the scuares is your method to use here

OpenStudy (anonymous):

First put it like this: (x^2+22x)+(y^2-14y)+49=0 now add a number inside paraentesis to make a perfect square. Remmeber also to substract that same number to keep equation unchanged: (x^2+22x +121)+(y^2-14y+49)+49-121-49=0 this way we get: (x+11)^2+(y-7)^2=121 do you recognize this equation?

OpenStudy (anonymous):

Yes i recognize :D

OpenStudy (anonymous):

ok then, :)